if a car can break at an acceleration of 20n/s2 what is the minimun distance in meters it needs to get to a full stop starting at a speed 44.7m /s
Answer #1

here,

the initial speed of car , u = 44.7 m/s

the final speed of car , v = 0 m/s^2

acceleration of car , a = - 20 m/s^2

let the stopping distance be s

using seccond equation of motion

v^2 - u^2 = 2 * a * s

0^2 - 44.7^2 = 2 * ( -20) * s

solving for s

s = 50 m

the stopping distance for car is 50 m